# Exploring the Dynamics of General Relativistic Binary-Single and Binary-Binary Encounters of Black Holes

### v1

@@ Abstract
In this proof-of-principle study, we demonstrate the capability of the numerical-relativity code BAM to simulate fully relativistic black-hole binary-single and binary-binary encounters that result in flybys, delayed or accelerated eccentric mergers, exchanges, and other more complex dynamical interactions of initially non-spinning, equal-mass black holes. Our results show that we are able to simulate the time evolution of $N$ black holes in relativistic scattering scenarios, which exhibit interesting dynamics and gravitational-wave signals. The dynamics of these systems show noticeable differences compared to analogous systems in post-Newtonian approximations up to 2.5PN. A key result is that the gravitational waveforms exhibit remarkable features that could potentially make them distinguishable from regular binary mergers.
@@ I Introduction
The $N$-body problem in full general relativity requires additional developments. After the 2005 breakthrough in numerical relativity, which for the first time made simulations of the late inspiral and merger of two black holes possible [36, 37, 38, 39], several fully relativistic studies of systems containing more than two black holes have also been carried out. These studies have either looked at arbitrary $N$-black-hole configurations [40, 41] (recently [42] evolved a cluster of 25 black holes), or they focused on the dynamics of systems containing a black hole orbiting in close proximity to a black-hole binary [43, 44, 45], but more general binary-single or binary-binary encounters have so far not been investigated in detail.
Since small changes in the initial conditions can lead to vastly different outcomes, it often does not suffice to look at only a few similar scattering configurations. Most of the major results for $N$-body scattering processes are obtained by running a large number of simulations (often more than hundreds of thousands or millions) and investigating the statistical properties of the outcomes. This poses a major challenge for numerical-relativity simulations as they are computationally expensive, such that it is unfeasible to produce even on the order of a thousand of such scattering experiments with reasonable time and energy resources.
@@
The paper is structured as follows. In Section II we provide a quick summary of how we compute the initial data and how these are evolved in time, including the numerical methods we use to solve the corresponding equations. In Section III we explain how we set up our binary-single and binary-binary scattering experiments and what needs to be taken into consideration in order to obtain accurate results. Our first simulations are presented in Section IV, where we show examples of different outcomes of the encounters, their dynamics, as well as their gravitational-wave signals. A summary of our results can be found in Section V. Throughout this paper we use geometric units where $G=c=1$.
@@ II Methods
Numerical relativity offers a variety of methods for the solution of the Einstein equations (e.g. [46, 47, 48, 49]) and here we introduce a specific choice suitable for more than two black holes in vacuum. The basic framework is the standard 3+1 decomposition. For $N>2$, the puncture method is a very flexible choice since it leads to minimal complications, say, compared to black hole excision methods. In principle, an arbitrary number of black hole punctures can move freely across a numerical grid without requiring special techniques for adapted coordinates or excision regions. This section introduces the relevant aspects of the moving puncture method for this work.
@@ II.1 Initial Data
where $\bar{A}_{ij}$ is trace-free. We further adopt Cartesian coordinates, assume conformal flatness ($\tilde{\gamma}_{ij}=\eta_{ij}$) and maximal slicing ($K=0$) which greatly simplifies and decouples the constraint equations for $\psi$ and $\bar{A}_{ij}$. With these choices, the momentum constraint (1) takes the form
@@
Bowen and York [51] found a non-trivial analytic solution to the momentum constraint, which describes $N$ black holes, where the $\alpha$-th black hole has an ADM linear momentum $P_{\alpha}^{i}$ and angular momentum $J_{\alpha}^{i}$, given by {dmath} ¯A^ij = ∑_α=1^N [32rCα2 ( P_α^i l_C_α^j + P_α^j l_C_α^i - (η^ij - l_C_α^i l_C_α^j) l_C_α^k P_α,k ) + 6rCα3 l_C_α^(i ϵ^j)kl J_α,k l_C_α, l ]. Here $r_{C_{\alpha}}=|x^{i}-C_{\alpha}^{i}|$ and $l_{C_{\alpha}}^{i}=(x^{i}-C_{\alpha}^{i})/r_{C_{\alpha}}$, where $C_{\alpha}^{i}$ are the coordinates of the center of the $\alpha$-th black hole.
@@ II.2 Time Evolution
$\begin{split}(\partial_{t}-\mathcal{L}_{\beta})\tilde{A}_{ij}={}&\chi[-\tilde{D}_{i}\tilde{D}_{j}\alpha+\alpha R_{ij}]^{TF}\\ &+\alpha\left[K\tilde{A}_{ij}-2\tilde{A}_{ik}\tilde{A}^{k}{}_{j}\right],\end{split}$ (14)
@@ II.4 Numerical Methods
With regard to loops over the $N$ black holes, some parts of the code require nested double loops. For the moving box algorithm using bounding boxes, we have to check whether any pair of boxes overlap, but apart from parallelization, this is straightforward. A similar consideration applies to the apparent-horizon finding. While marginally trapped surfaces can be found straightforwardly near each individual puncture, we have not yet considered the issue of finding such surfaces surrounding two or more punctures simultaneously. In principle, BAM’s flow algorithm can handle these situations by starting with appropriate initial surfaces, but this requires additional coding.
The gravitational wave extraction is performed by using the Newman-Penrose scalar $\Psi_{4}$ as described in [53]. A fundamental issue arises for configurations with more than two black holes. We typically require a reference frame with a center and an axis to define a tetrad basis. There are well-known issues related to the choice of frame. Here the issue is not that we have to loop over $N$ sources; in fact, there is just one extraction sphere, but there are different choices for this sphere, leading to various mode mixing effects.
@@ II.5 Post-Newtonian Methods
For the post-Newtonian (PN) approximations of the binary-single encounters up to 2.5PN we use the three-body Hamiltonian and the equations of motion presented in [66, 67, 68, 69, 70] and we performed similar tests to verify the validity of our results. We integrate the equations of motion using a 5th order Runge-Kutta scheme with an adaptive step size to keep the relative error below $10^{-6}$.
@@ III Configuration Setup
Figure 1 depicts the setup schematically. Both the black-hole binary’s center of mass and the single black hole (or the other black-hole binary’s center of mass) start on the $x$-axis at locations ±$\pm X$, resulting in an initial separation of $2X$. The momenta ±$\pm\vec{p}$ lie in the $xy$-plane and have components
±$\pm\vec{p}=\pm|\vec{p}|\begin{pmatrix}-\sqrt{1-\left(\frac{b}{2X}\right)^{2}}\\ \frac{b}{2X}\\ 0\end{pmatrix},$ (21)
@@
We use three different grid setups, depending on the required size of the domain in which the black holes should be able to move. In all of our simulations, the grid hierarchy contains 11 levels indexed by $L=0,1,...,10$. Levels 0 to 5 consist of fixed cubical boxes with $n$ grid points in each direction and a grid spacing of $h_{L}=h_{0}/2^{L}$, that are centered at the origin. Levels 6 to 10 consist of moving cubical boxes that are centered at the locations of the punctures, with $n/2$ grid points in each direction and a grid spacing of $h_{L}=h_{0}/2^{L}$. Our values for $n$ and $h_{0}$ for the three different grid setups can be found in Tables 1 to 3. In some cases, the number of grid points is slightly adjusted to guarantee a total amount of $24n+$ 2 grid points in each direction.
@@ IV.1 Weak Binary-Single Scattering (BSS-01)
Our first case study is a black-hole binary-single scattering with $X=100$ M (BSS-01a) and an impact parameter $b=50$ M that is several times greater than the initial coordinate separation of the black holes that make up the binary, leading to a relatively weak perturbation of the binary dynamics. The binary is oriented such that the normal vector of the orbital plane is parallel to the momentum vector $-\vec{p}$, i.e. the orbital plane of the binary is chosen to be approximately perpendicular to the scattering motion occurring mostly in the $xy$-plane. The initial relative approach velocity of the binary and the single black hole is 0.223. A detailed list of all the initial parameters can be found in Table 4 in Appendix A.
@@
Figure 8: Top left: Puncture trajectories of the black holes of BSS-02 (colored blue for the black holes in the binary and orange for the single black hole). The colored crosses correspond to the times marked by the dashed lines in the other plots. Top right: Gravitational-wave luminosity extracted at the radius $R=400$ M and on the level $L=2$. Bottom: The real part (blue), the imaginary part (orange), and the absolute value (green) of the 22-mode of $R\Psi_{4}$, extracted at the same radius and level.
@@ IV.2 Complex Interaction With a Double Merger After a Strong Binary-Single Encounter (BSS-02)
Figure 9: The $zy$-plane view on three-dimensional volume rendering of the amplitude of $|R\Psi_{4}|$ of BSS-02 at different times and limited to the emission sphere of radius $R=200$. The data is shown only at the isosurfaces of selected values of the amplitude to highlight the emission features, i.e., other amplitude values also contain emission. The puncture trajectories of the black holes are indicated in white. The times are picked to capture the emission from the moving binary ($t=996$), the first two bursts of emission after encounter with the incoming black hole ($t=1226$), and the last burst of emission from the coalescence of the newly formed binary ($t=1952$).
@@
Figure 8 shows the puncture trajectories of the three black holes (top left), the gravitational-wave luminosity (top right), and the 22-mode of $R\Psi_{4}$, extracted at the radius $R=400$ M and on the level $L=2$. The plot of the gravitational-wave luminosity shows four distinct peaks, which we marked with differently colored dashed lines and which correspond to the dashed lines in the plot for $R\Psi_{4,22}$. One can connect the peaks in the gravitational-wave emission to specific events in the binary-single encounter. We illustrate this connection by adding crosses to the puncture tracks at the retarded times $t_{R}=t-R$, which approximately correspond to the moments in time $t$ marked by the dashed lines. The fact that the interactions do not happen at the origin of the grid and the space is not flat makes this a rough estimate, but it should allow for a qualitative assessment. The first peak (colored red) is connected with the first strong encounter, where the single black hole pulls both black holes in the binary in a different direction. The second peak (colored magenta) is associated with the second interaction of the single black hole, primarily with one of the binary members, which temporarily ejects it from the strong interaction region. The third peak (colored turquoise) is created at the merger of the single black hole with one of the binary members, and the fourth peak (colored brown) corresponds to the merger of the remaining black holes.
@@
We now discuss the gravitational-wave emission and extraction in more detail for this example. The 22-mode of $R\Psi_{4}$ starts with a regular gravitational-wave signal of a quasi-circular binary and then shows several peaks that are associated with the peaks in the gravitational-wave luminosity. The first two peaks cannot be distinguished in the 22-mode, and all visible peaks show high-frequency oscillations in the absolute value of $R\Psi_{4,22}$. We suggest that this is not a physical effect but rather a result of the source being displaced from the origin of the grid as well as moving, rotating, and accelerating within the extraction sphere, which causes mode mixing [75, 76, 77], anisotropic signal arrival times on the extraction sphere across different directions, blueshifting of the signals in the direction of motion, and redshifting in the opposite direction. For comparison, we provide the 22-mode and the 44-mode of a merger signal of a black-hole binary that is moving uniformly with respect to the origin of the coordinate grid in Figure 20 in Appendix A. The gravitational-wave luminosity, which is computed by using the total $\Psi_{4}$, which contains all the modes, does not show these spurious oscillations. Furthermore, the emission of gravitational waves in $N$-body systems can originate from very different locations and tends to be highly anisotropic, which is illustrated in Figure 9. The emissions from the initial binary are primarily in the $z$-direction (see the image on the left). In contrast, the first three bursts are sent in distinct directions that are mostly perpendicular to the $z$-axis. The emission of the first two bursts is depicted in the center panel of Figure 9. The last burst is primarily sent in the positive and negative $z$-direction (see the image on the right). Because of this and the mode mixing, the 22-mode is not necessarily the most dominant and representative mode. As an alternative, one can reconstruct $R\Psi_{4}$ on the extraction sphere using all of the available modes and evaluate it in different directions. This is illustrated in Figure 10, which shows $R\Psi_{4}$ reconstructed from all modes from $l=2$ to $l=4$ for four different directions. The directions were picked such that each direction maximizes the amplitude of the absolute value of $R\Psi_{4}$ for a given peak. The spurious oscillations in the absolute value of $R\Psi_{4}$ are still present but slightly weaker than in the 22-mode. Due to the mode mixing, modes with higher $l$ also need to be considered and are no longer subdominant. We tried using all the modes up to $l=6$ which did not significantly reduce the oscillations but slightly changed the amplitudes of some peaks. When evaluating the full $\Psi_{4}$ computed at each grid point in the simulation across various locations on the extraction sphere, oscillations in $|R\Psi_{4}|$ persist, suggesting that mode mixing and considering only a finite number of modes is not the only contributor to these oscillations. We expect that extracting the signal at much larger extraction radii and extending the signal to null-infinity can alleviate some of these issues. Despite that, one can clearly see, both from Figure 9 and Figure 10, that each gravitational-wave burst is sent in very different directions. For other directions, the emission of a burst might be very small, or two peaks that are close to each other might become indistinguishable. For the systems that we present in the following sections, we pick a direction that shows all the existing features of the gravitational-wave signal most clearly.
@@
Figure 10: The real part (blue), the imaginary part (orange), and the absolute value (green) of $R\Psi_{4}$ for the system BSS-02 extracted at the radius $R=400$ M and on the level $L=2$. $R\Psi_{4}$ is reconstructed from all modes from $l=2$ to $l=4$ and is evaluated in different directions (indicated at the top left). The colored dashed lines correspond to those in Figure 8.

### v2

@@ Abstract
In this exploratory study, we demonstrate the capability of the numerical-relativity code BAM to simulate fully relativistic black-hole binary-single and binary-binary encounters. While previous work has demonstrated the general capability of numerical-relativity frameworks to evolve spacetimes with $N$ black holes, detailed explorations of such encounters remain limited. We focus on scenarios involving initially non-spinning, equal-mass black holes that result in a variety of dynamical outcomes, including flybys, delayed or accelerated eccentric mergers, exchanges, and more complex interactions. Our results show that we can reliably simulate scattering experiments involving three and four black holes, which exhibit interesting dynamics and gravitational-wave signals. The dynamics of these systems show noticeable differences compared to analogous systems in post-Newtonian approximations up to 2.5PN. A key result is that the gravitational waveforms exhibit remarkable features that could potentially make them distinguishable from regular binary mergers.
@@ I Introduction
In full general relativity, prior to the $N$-body problem, even the 2-body problem required substantial developments. Early examples for non-axisymmetric black-hole binaries using the “puncture method” include grazing collisions [7] and the first complete binary orbit before merger [25]. The year 2005 marks a breakthrough for the late inspiral and merger of two black holes [76, 28, 13], where either dynamical black-hole excision or a “moving puncture” approach was used. Such methods generalize directly to $N$ black holes, as first demonstrated for moving punctures in [65]. After that, several fully relativistic studies of systems containing more than two black holes have been carried out, but actually not that many. These studies have either looked at arbitrary $N$-black-hole configurations [65, 46] (recently [14] evolved a cluster of 25 black holes), or they focused on the dynamics of systems containing a black hole orbiting in close proximity to a black-hole binary [29, 40, 41], but more general binary-single or binary-binary encounters have so far not been investigated in detail.
Since small changes in the initial conditions can lead to vastly different outcomes, it often does not suffice to look at only a few similar scattering configurations. Most of the major results for $N$-body scattering processes are obtained by running a large number of simulations (often more than hundreds of thousands or millions) and investigating the statistical properties of the outcomes. This poses a major challenge for numerical-relativity simulations as they are computationally expensive, such that it is unfeasible to produce even on the order of a thousand of such scattering experiments within a reasonable time.
@@
The paper is structured as follows. In Section II we provide a quick summary of how we compute the initial data and how these are evolved in time, including the numerical methods we use to solve the corresponding equations. In Section III we explain how we set up our binary-single and binary-binary scattering experiments and what needs to be taken into consideration in order to obtain accurate results. Our first simulations are presented in Section IV, where we show examples of different outcomes of the encounters, their dynamics, as well as their gravitational-wave signals. A summary of our results can be found in Section V. In this paper we use metric signature (,+,+,+)(-,+,+,+), geometric units where $G=c=1$, and a mass scale $M$ which in code units is $M=1$.
@@ II Methods
Numerical relativity offers a variety of methods for the solution of the Einstein equations (e.g. [17, 8, 15, 50, 27]). Here, we adopt a specific approach suitable for vacuum spacetimes with N black holes, based on the standard 3+1 decomposition. For $N>2$, the puncture method provides a flexible framework: black hole punctures can move freely across a numerical grid without requiring adapted coordinates or excision. This section introduces the relevant aspects of the moving puncture method for this work.
@@ II.1 Initial Data
where $\bar{A}_{ij}$ is trace-free. We further adopt Cartesian coordinates, assume conformal flatness ($\tilde{\gamma}_{ij}=\eta_{ij}$) and maximal slicing ($K=0$), which greatly simplifies and decouples the constraint equations for $\psi$ and $\bar{A}_{ij}$. With these choices, the momentum constraint (1) takes the form
@@
Bowen and York [19] found a non-trivial analytic solution to the momentum constraint, which describes $N$ black holes, where the $\alpha$-th black hole has an ADM linear momentum $P_{\alpha}^{i}$ and angular momentum $J_{\alpha}^{i}$, given by
$\bar{A}^{ij}=\sum_{\alpha=1}^{N}\left[\frac{3}{2r_{C_{\alpha}}^{2}}\left(P_{\alpha}^{i}l_{C_{\alpha}}^{j}+P_{\alpha}^{j}l_{C_{\alpha}}^{i}-(\eta^{ij}-l_{C_{\alpha}}^{i}l_{C_{\alpha}}^{j})l_{C_{\alpha}}^{k}P_{\alpha,k}\right)+\frac{6}{r_{C_{\alpha}}^{3}}l_{C_{\alpha}}^{(i}\epsilon^{j)kl}J_{\alpha,k}l_{C_{\alpha},l}\right].$ (6)
Here $r_{C_{\alpha}}=|x^{i}-C_{\alpha}^{i}|$ and $l_{C_{\alpha}}^{i}=(x^{i}-C_{\alpha}^{i})/r_{C_{\alpha}}$, where $C_{\alpha}^{i}$ are the coordinates of the center of the $\alpha$-th black hole.
@@ II.2 Time Evolution
[…]
@@ II.4 Numerical Methods
The gravitational wave extraction is performed by using the Newman-Penrose scalar $\Psi_{4}$ as described in [24]. A fundamental issue arises for configurations with more than two black holes. We typically require a reference frame with a center and an axis to define a tetrad basis. There are well-known issues related to the choice of frame, leading to various effects such as mode mixing.
The accuracy of our numerical simulations is evaluated in Appendix B. In particular the constraint violations, the conservation of mass, linear momentum and angular momentum as well as the behavior of the gravitational-wave signals at different extraction radii are discussed.
@@ II.5 Post-Newtonian Methods
For the post-Newtonian (PN) approximations of the binary-single encounters up to 2.5PN we use the three-body Hamiltonian and the equations of motion presented in [85, 64, 47, 48, 18, 93] and we performed similar tests to verify the validity of our results. We integrate the equations of motion using a 5th order Runge-Kutta scheme with an adaptive step size to keep the relative error below $10^{-6}$.
@@ III Configuration Setup
Figure 1 depicts the setup schematically. Both the black-hole binary’s center of mass and the single black hole (or the other black-hole binary’s center of mass) start on the $x$-axis at locations $\pm X$, resulting in an initial separation of $2X$. The momenta $\pm\vec{p}$ lie in the $xy$-plane and have components
$\pm\vec{p}=\pm|\vec{p}|\begin{pmatrix}-\sqrt{1-\left(\frac{b}{2X}\right)^{2}}\\ \frac{b}{2X}\\ 0\end{pmatrix}$, (22)
@@
We use three different grid setups, depending on the required size of the domain in which the black holes should be able to move. In all of our simulations, the grid hierarchy contains 11 levels indexed by L=0,1,,10L=0,1,...,10. Levels 0 to 5 consist of fixed cubical boxes with $n$ grid points in each direction and a grid spacing of $h_{L}=h_{0}/2^{L}$, that are centered at the origin. Levels 6 to 10 consist of moving cubical boxes that are centered at the locations of the punctures, with $n/2$ grid points in each direction and a grid spacing of $h_{L}=h_{0}/2^{L}$. Our values for $n$ and $h_{0}$ for the three different grid setups can be found in Tables 1 to 3. In some cases, the number of grid points is slightly adjusted to guarantee a total amount of $4n+2$ grid points in each direction.
@@ IV.1 Weak Binary-Single Scattering (BSS-01)
Our first case study is a black-hole binary-single scattering with $X=100M$ (BSS-01a) and an impact parameter $b=50M$ that is several times greater than the initial coordinate separation of the black holes that make up the binary, leading to a relatively weak perturbation of the binary dynamics. The binary is oriented such that the normal vector of the orbital plane is parallel to the momentum vector $\vec{p}$, i.e. the orbital plane of the binary is chosen to be approximately perpendicular to the scattering motion occurring mostly in the $xy$-plane. The initial relative approach velocity of the binary and the single black hole is 0.223. A detailed list of all the initial parameters can be found in Table 4 in Appendix A.
@@
Figure 8: Top left: Puncture trajectories of the black holes of BSS-02 (colored blue for the black holes in the binary and orange for the single black hole). The colored crosses correspond to the times marked by the dashed lines in the other plots. Top right: Gravitational-wave luminosity extracted at the radius $R=400M$ and on the level $L=2$. Bottom: The real part (blue), the imaginary part (orange), and the absolute value (green) of the 22-mode of $RM\Psi_{4}$, extracted at the same radius and level.
@@ IV.2 Complex Interaction With a Double Merger After a Strong Binary-Single Encounter (BSS-02)
Figure 9: The zyzy-plane view on three-dimensional volume rendering of the amplitude of $|RM\Psi_{4}|$ of BSS-02 at different times and limited to the emission sphere of radius $R=200M$. The data is shown only at the isosurfaces of selected values of the amplitude to highlight the emission features, i.e., other amplitude values also contain emission. The puncture trajectories of the black holes are indicated in white. The times are picked to capture the emission from the moving binary ($t=996M$), the first two bursts of emission after encounter with the incoming black hole ($t=1226M$), and the last burst of emission from the coalescence of the newly formed binary ($t=1952M$).
@@
Figure 8 shows the puncture trajectories of the three black holes (top left), the gravitational-wave luminosity (top right), and the 22-mode of $RM\Psi_{4}$, extracted at the radius $R=400M$ and on the level $L=2$. The plot of the gravitational-wave luminosity shows four distinct peaks, which we marked with differently colored dashed lines and which correspond to the dashed lines in the plot for $RM\Psi_{4,22}$. One can connect the peaks in the gravitational-wave emission to specific events in the binary-single encounter. We illustrate this connection by adding crosses to the puncture tracks at the retarded times $t_{R}=t-R$, which approximately correspond to the moments in time $t$ marked by the dashed lines. The fact that the interactions do not happen at the origin of the grid and the space is not flat makes this a rough estimate, but it should allow for a qualitative assessment. The first peak (colored red) is connected with the first strong encounter, where the single black hole pulls both black holes in the binary in a different direction. The second peak (colored magenta) is associated with the second interaction of the single black hole, primarily with one of the binary members, which temporarily ejects it from the strong interaction region. The third peak (colored turquoise) is created at the merger of the single black hole with one of the binary members, and the fourth peak (colored brown) corresponds to the merger of the remaining black holes.
@@
We now discuss the gravitational-wave emission and extraction in more detail for this example. The 22-mode of $RM\Psi_{4}$ starts with a regular gravitational-wave signal of a quasi-circular binary and then shows several peaks that are associated with the peaks in the gravitational-wave luminosity. The first two peaks cannot be distinguished in the 22-mode, and all visible peaks show high-frequency oscillations in the absolute value of $RM\Psi_{4,22}$. We suggest that this is not a physical effect but rather a result of the source being displaced from the origin of the grid as well as moving, rotating, and accelerating within the extraction sphere, which causes mode mixing [20, 30, 92], anisotropic signal arrival times on the extraction sphere across different directions, blueshifting of the signals in the direction of motion, and redshifting in the opposite direction. For comparison, we provide the 22-mode and the 44-mode of a merger signal of a black-hole binary that is moving uniformly with respect to the origin of the coordinate grid in Figure 20 in Appendix A. The gravitational-wave luminosity, which is computed by using the total $\Psi_{4}$, which contains all the modes, does not show these spurious oscillations. Furthermore, the emission of gravitational waves in $N$-body systems can originate from very different locations and tends to be highly anisotropic, which is illustrated in Figure 9. The emissions from the initial binary are primarily in the $z$-direction (see the image on the left). In contrast, the first three bursts are sent in distinct directions that are mostly perpendicular to the $z$-axis. The emission of the first two bursts is depicted in the center panel of Figure 9. The last burst is primarily sent in the positive and negative $z$-direction (see the image on the right). Because of this and the mode mixing, the 22-mode is not necessarily the most dominant and representative mode. As an alternative, one can reconstruct $RM\Psi_{4}$ on the extraction sphere using all of the available modes and evaluate it in different directions. This is illustrated in Figure 10, which shows $RM\Psi_{4}$ reconstructed from all modes from $l=2$ to $l=4$ for four different directions. The directions were picked such that each direction maximizes the amplitude of the absolute value of $RM\Psi_{4}$ for a given peak. The spurious oscillations in the absolute value are still present but slightly weaker than in the 22-mode. Due to the mode mixing, modes with higher $l$ also need to be considered and are no longer subdominant. We tried using all the modes up to $l=6$ which did not significantly reduce the oscillations but slightly changed the amplitudes of some peaks. When evaluating the full $\Psi_{4}$ computed at each grid point in the simulation across various locations on the extraction sphere, oscillations in $|RM\Psi_{4}|$ persist, suggesting that mode mixing and considering only a finite number of modes is not the only contributor to these oscillations. We expect that extracting the signal at much larger extraction radii and extending the signal to null-infinity can alleviate some of these issues. Despite that, one can clearly see, both from Figure 9 and Figure 10, that each gravitational-wave burst is sent in very different directions. For other directions, the emission of a burst might be very small, or two peaks that are close to each other might become indistinguishable. For the systems that we present in the following sections, we pick a direction that shows all the existing features of the gravitational-wave signal most clearly.
@@
Figure 10: The real part (blue), the imaginary part (orange), and the absolute value (green) of $RM\Psi_{4}$ for the system BSS-02 extracted at the radius $R=400M$ and on the level $L=2$. $RM\Psi_{4}$ is reconstructed from all modes from $l=2$ to $l=4$ and is evaluated in different directions (indicated at the top left). The colored dashed lines correspond to those in Figure 8.